OpenStudy (anonymous):

Solve the equation. Show your work. 45 = 3b + 69

OpenStudy (anonymous):

45=3b+69 45-69=3b+69-68 -24=3b b=-8

OpenStudy (anonymous):

second sentence correction 45-69=3b+69-69

OpenStudy (anonymous):

45-69=-24 -24/3=-8